1. Using the given information answer the following questions with corresponding answers.

(a)

Tryptophan is a diprotic acid (Ka1 = 1.48 LaTeX: \times× 10-3 and Ka2 = 4.07 LaTeX: \times× 10-10) that can have three different form in solution: H2T+, HT, and T-. What would be the pH of the solution containing 0.045M HT?

Group of answer choices

6.86

6.12

7.43

12.23

7.06

(b)

Tryptophan is a diprotic acid (Ka1 = 1.48 LaTeX: \times× 10-3 and Ka2 = 4.07 LaTeX: \times× 10-10) that can have three different form in solution: H2T+, HT, and T-. What would be the concentration of H2T+ in the solution of 0.045M HT?

Group of answer choices

4.2 LaTeX: \times× 10-5 M

0.80 LaTeX: \times× 10-5 M

5.1 LaTeX: \times× 10-5 M

2.3 LaTeX: \times× 10-5 M

3.7 LaTeX: \times× 10-5 M

(c)

Tryptophan is a diprotic acid (Ka1 = 1.48 LaTeX: \times× 10-3 and Ka2 = 4.07 LaTeX: \times× 10-10) that can have three different form in solution: H2T+, HT, and T-. What would be the concentration of A- in the solution of 0.045M HT?

Group of answer choices

0.38 LaTeX: \times× 10-5 M

3.3 LaTeX: \times× 10-5 M

1.3 LaTeX: \times× 10-5 M

2.4 LaTeX: \times× 10-5 M

4.7 LaTeX: \times× 10-5 M

(d)

Serine is a diprotic acid (Ka1 = 6.17 LaTeX: \times× 10-3 and Ka2 = 7.08 LaTeX: \times× 10-10) that can have three different forms in solution: H2S+, HS, and S- (Note: S is not sulfur in these formulas). What would be the pH of the solution containing 0.035M Na+S-?

Group of answer choices

9.88

10.23

10.85

9.02

3.15
Answer #1

Tryptophan ihon - HI sincee dit act as aeed amphiprotic so - teigika, & {- leg Kaz} = -deg (1.60*10) + {-lea, 9-03 Xpo} 20831:49 X 10-3 = 0.045 % 10 [1₂7] 0.045 x 106-11 1.48X10-8. [h, T = 1 = prog5 x 7680 - = 0.045 xlvo X10 1.488 0.0304 x 10 [7] =